# Secondary terms in the distribution of genus numbers of cubic fields

### v1

@@ 1 Introduction
Our proof largely follows the strategy of McGown and Tucker [MT23], but we introduce a distinct method to handle the specific condition at the prime $3$. To establish the main theorems, it is necessary to count cubic fields $F$ whose fundamental discriminants $d_{F}$ satisfy a congruence condition $d_{F}\equiv 1\pmod{3}$ (see (14)). While McGown and Tucker addressed this using the geometry of numbers, we employ Dirichlet characters modulo $3$. The counting of cubic fields twisted by Dirichlet characters was carried out by Taniguchi and Thorne [TT13]. The proof of the improved uniform estimates of [BTT24] are directly applicable to this twisted count, yielding error terms of the same quality as those for the untwisted count. This allows us to obtain the improved error estimates in our main theorems. As far as we are aware, this is the first time the twisted counting functions in [TT13] have been applied to further arithmetic statistics.
@@
The organization of this paper is as follows: In Section 2, we summarize the necessary results on the density of cubic fields. In particular, we state the uniform error term estimates for counting cubic fields with prescribed local specifications, which are slight generalizations of the results in [BTT24]. In Section 3, we provide the proofs of our main results, Theorems 1.3 and 1.4. We first recall Fröhlich’s formula, which shows that the genus number is determined by local properties. Then, using the discriminant decomposition we analyze the counting functions for cubic fields with fixed squarefree parts of the discriminant. By summing these contributions, we derive the asymptotic formulas for the distribution of genus numbers.
@@ 2 Preliminaries on Cubic Field Densities
In studying the distribution of cubic fields satisfying given local conditions, the following two theorems (Theorem 2.1, and Theorem 2.2) serve as our main tools. They improve the modulo $1$ or $3$ cases of Theorem 6.2 in [TT13] by incorporating the error estimates of [BTT24]. Specifically, Theorems 2.1 and 2.2 generalize Theorems 1.3 and 1.4 in [BTT24], respectively. Exactly the same proofs work.
Theorem 2.1.
Let $\Sigma=\left(\Sigma_{v}\right)_{v}$ denote an almost ordinary collection of local specifications. Let $\chi$ be a primitive Dirichlet character modulo $1$ or $3$; that is, $\chi$ is either the trivial character $1\pmod{1}$ or the Legendre character $\chi_{3}\pmod{3}$. We define
$N\left(X,\Sigma,\chi\right)\coloneqq\sum_{\begin{subarray}{c}\left|\operatorname{Disc}\left(F\right)\right|\leq X\\ F\in\Sigma\end{subarray}}\chi\left(\frac{\operatorname{Disc}\left(F\right)}{\operatorname{Disc}_{3}\left(F\right)}\right),$ (7)
where $\operatorname{Disc}_{3}\left(F\right)$ denote the $3$-part of the discriminant of $F$. Then, we have
$N\left(X,\Sigma,\chi\right)=\begin{dcases}C\left(\Sigma\right)\frac{1}{12\zeta\left(3\right)}X+K\left(\Sigma\right)\frac{4\zeta\left(1/3\right)}{5{\Gamma\left(2/3\right)}^{3}\zeta\left(5/3\right)}X^{5/6}+O_{\varepsilon}\left(E\left(\Sigma\right)X^{2/3+\varepsilon}\right)&\text{if }\chi=1;\\ O_{\varepsilon}\left(E\left(\Sigma\right)X^{2/3+\varepsilon}\right)&\text{if }\chi=\chi_{3},\end{dcases}$ (8)
@@
where $v$ runs over all places of $\mathbb{Q}$. The local factors $C_{\infty}\left(\Sigma_{\infty}\right)$, $K_{\infty}\left(\Sigma_{\infty}\right)$, and $E_{\infty}\left(\Sigma_{\infty}\right)$ are defined by $C_{\infty}\left(\pm\right)\coloneqq C^{\pm}$, $K_{\infty}\left(\pm\right)\coloneqq K^{\pm}$, and $E_{\infty}\left(\Sigma_{\infty}\right)\coloneqq 1$. For each prime $p$, the local factors $C_{p}\left(\Sigma_{p}\right)$, $K_{p}\left(\Sigma_{p}\right)$, and $E_{p}\left(\Sigma_{p}\right)$ are defined as in Theorem 1.3 of [BTT24].
@@
Theorem 2.2.
For each collection of cubic local specifications $\Sigma=\left(\Sigma_{v}\right)_{v}$ and $\chi\in\left\{1,\chi_{3}\right\}$, let $E\left(X,\Sigma,\chi\right)$ denote the error term in estimating $N\left(X,\Sigma,\chi\right)$ in Theorem 2.1. Then for any $\varepsilon>0$, we have
$\sum_{r\leq R}\sum_{s}\left|E\left(X_{r,s},\Sigma^{r,s},\chi\right)\right|\ll_{\varepsilon}X^{2/3+\varepsilon}R^{2/3}\prod_{p\mid U}E_{p}\left(\Sigma_{p}\right)$ (10)
@@ 3 Proofs of Main Results
The contribution of cyclic fields is negligible for the main results of this paper. First, regarding Theorem 1.3, the number of cyclic cubic fields with discriminant up to $X$ is known to be $O\left(X^{1/2}\right)$ (see [MT23, Section 2]). Second, regarding Theorem 1.4, we verify that the contribution of cyclic fields to the moments is negligible. Recall that for a squarefree positive integer $f$ coprime to $3$, the number of cyclic cubic fields $F$ satisfying $f_{F}=f$ is at most $2^{\omega\left(f\right)}$, and their genus numbers satisfy $g_{F}\leq 3^{\omega\left(f\right)}$ (see [MT23, Section 4]). Under the assumption $\operatorname{Re}\left(z\right)\leq G\left(X\right)$ with $G\left(X\right)=o\left(\log\log X\right)$, Théorème 11 of [Rob83] implies that for any $\varepsilon>0$, there exists a constant $C^{\prime}_{G,\varepsilon}$ depending on $G$ and $\varepsilon$ such that for all $X>C^{\prime}_{G,\varepsilon}$, we have
@@
Then we have $M\left(X,f\right)=N\left(X,\Sigma^{f},1\right)$, where $1$ denote the trivial Dirichlet character modulo $1$. For convenience, we define
@@
$\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E\left(X,f\right)\right|=\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E\left(X,\Sigma^{f},1\right)\right|\ll_{\varepsilon}X^{2/3+\varepsilon}\prod_{p\mid 3c\left(\Sigma\right)t}E\left(\Sigma_{p}\right)\ll_{\varepsilon}{c\left(\Sigma\right)}^{2/3}X^{2/3+\varepsilon}.$ (22)
@@
Next, we consider $M^{\prime}\left(X,f\right)$. To handle the condition $d_{F}\equiv 1\pmod{3}$, we make use of the orthogonality of Dirichlet characters modulo $3$. For this purpose, we define
$\widetilde{\mathcal{F}^{\prime}}\left(f\right)\coloneqq\left\{F\in\mathcal{F}\left(f\right)\mid 3\text{ is totally ramified in }F\text{ and }3\nmid d_{F}\right\},$
and for $\chi\in\left\{1,\chi_{3}\right\}$, we define
$\widetilde{M^{\prime}}\left(X,f,\chi\right)\coloneqq\sum_{\begin{subarray}{c}\left|\operatorname{Disc}\left(F\right)\right|<X\\ F\in\widetilde{\mathcal{F}^{\prime}}\left(f\right)\end{subarray}}\chi\left(d_{F}\right).$
According to [TT14, Lemma 2.1], for a non-cyclic cubic field $F$, the condition that $3$ is totally ramified in $F$ and that $d_{F}$ is not divisible by $3$ is equivalent to the condition that the discriminant of $F$ has $3$-adic valuation exactly equal to $4$. Let $\mathscr{A}_{3}^{\left(4\right)}$ denote the subset of $\mathscr{A}_{3}$ consisting of cubic étale $\mathbb{Q}_{3}$-algebras corresponding to such fields.
@@
$\alpha^{\prime}\left(f\right)$ $\coloneqq\frac{\left(1+a_{3}\right)C_{\infty}\left(\Sigma_{\infty}\right)C_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)}{24\zeta\left(2\right)}a_{tf}\prod_{p\mid c\left(\Sigma\right)}C_{p}\left(\Sigma_{p}\right)\left(1+a_{p}\right),$
$\beta^{\prime}\left(f\right)$ $\coloneqq\frac{2\left(1+b_{3}\right)\zeta\left(1/3\right)K_{\infty}\left(\Sigma_{\infty}\right)K_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)}{5{\Gamma\left(2/3\right)}^{3}}b_{tf}\prod_{p}\left(1-\frac{p^{1/3}+1}{p\left(p+1\right)}\right)\prod_{p\mid c\left(\Sigma\right)}K_{p}\left(\Sigma_{p}\right)\left(1+b_{p}\right),$
@@
By the orthogonality of Dirichlet characters modulo $3$, we obtain
$M^{\prime}\left(X,f\right)=\sum_{\begin{subarray}{c}\left|\operatorname{Disc}\left(F\right)\right|<X\\ F\in\widetilde{\mathcal{F}^{\prime}}\left(f\right)\\ d_{F}\equiv 1\!\!\!\!\!\pmod{3}\end{subarray}}1=\sum_{\begin{subarray}{c}\left|\operatorname{Disc}\left(F\right)\right|<X\\ F\in\widetilde{\mathcal{F}^{\prime}}\left(f\right)\end{subarray}}\frac{1+\chi_{3}\left(d_{F}\right)}{2}.$
@@
$\Sigma^{3,f}_{v}\coloneqq\begin{dcases}\Sigma_{v}\cap\mathscr{T}_{v}&\text{if }v\mid tf;\\ \Sigma_{v}\cap\mathscr{A}^{\prime}_{v}&\text{if }v\nmid 3tf;\\ \Sigma_{v}&\text{if }v=\infty;\\ \Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}&\text{if }v=3,\end{dcases}$ (24)
then for $F\in\widetilde{\mathcal{F}^{\prime}}\left(f\right)$, $\chi_{3}\left(\frac{\operatorname{Disc}\left(F\right)}{\operatorname{Disc}_{3}\left(F\right)}\right)=\chi_{3}\left(d_{F}{f_{F}}^{2}\right)=\chi_{3}\left(d_{F}\right)$, and hence
$M^{\prime}\left(X,f\right)=\frac{1}{2}N\left(X,\Sigma^{3,f},1\right)+\frac{1}{2}N\left(X,\Sigma^{3,f},\chi_{3}\right),$
where $1$ denote the trivial Dirichlet character modulo $1$.
As for $N\left(X,\Sigma^{3,f},\chi_{3}\right)$, we have $N\left(X,\Sigma^{3,f},\chi_{3}\right)\ll_{\varepsilon}E\left(\Sigma^{3,f}\right)X^{2/3+\varepsilon}$. Regarding $N\left(X,\Sigma^{3,f},1\right)$, for all $v\neq 3$, we have $\Sigma^{3,f}_{v}=\Sigma^{f}_{v}$. Therefore, as in the previous proposition, $C\left(\Sigma^{3,f}\right)$, $K\left(\Sigma^{3,f}\right)$, and $E\left(\Sigma^{3,f}\right)$ are computed accordingly, and the proof of the first part is complete.
@@
$\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E^{\prime}\left(X,f\right)\right|=\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E\left(X,\Sigma^{3,f},1\right)\right|+\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E\left(X,\Sigma^{3,f},\chi_{3}\right)\right|\ll_{\varepsilon}c\left(\Sigma\right)^{2/3}X^{2/3+\varepsilon}.$ (25)
@@
For the explicit description of the algebras in $\mathscr{A}_{3}^{\left(4\right)}$ and the computation of the constants $C_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)$ and $K_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)$, we refer to Section 6.2 in [TT13].
@@
$I_{3}\left(k,\Sigma\right)$ $=\left(C_{3}\left(\Sigma_{3}\right)-\frac{C_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)}{2}\right)A_{k-l}\left(\Sigma\right)+\frac{C_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)}{2}A_{k-l-1}\left(\Sigma\right),$
$J_{3}\left(k,\Sigma\right)$ $=\left(K_{3}\left(\Sigma_{3}\right)-\frac{K_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)}{2}\right)B_{k-l}\left(\Sigma\right)+\frac{K_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)}{2}B_{k-l-1}\left(\Sigma\right).$
@@
The last equality is justified because $M\left(X,f\right)=M^{\prime}\left(X,f\right)=0$ for $f\geq\sqrt{X}$. We introduce this truncation to bound the contribution of the wights $3^{\psi\left(f\right)z}$ in the tail of the series in the above derivation. Defining $\alpha^{\Sigma}\left(z\right)$ and $\beta^{\Sigma}\left(z\right)$ as
@@
$\alpha^{\Sigma}\left(z\right)$ $\coloneqq\begin{aligned} \frac{3^{lz}\cdot C_{\infty}\left(\Sigma_{\infty}\right)}{12\zeta\left(2\right)}a_{t}&\left(C_{3}\left(\Sigma_{3}\right)+\frac{\left(3^{z}-1\right)}{2}C_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)\right)\\ &\quad\quad\quad\quad\quad\times\prod_{p\mid c\left(\Sigma\right)}C_{p}\left(\Sigma_{p}\right)\left(1+a_{p}\right)\prod_{p\in\mathcal{P}\cup\left\{3\right\}}\left(1+{n_{p}}^{z}a_{p}\right),\end{aligned}$ (29)
$\beta^{\Sigma}\left(z\right)$ $\coloneqq\begin{aligned} \frac{3^{lz}\cdot 4\zeta\left(1/3\right)K_{\infty}\left(\Sigma_{\infty}\right)}{5{\Gamma\left(2/3\right)}^{3}}b_{t}&\prod_{p}\left(1-\frac{p^{1/3}+1}{p\left(p+1\right)}\right)\left(K_{3}\left(\Sigma_{3}\right)+\frac{\left(3^{z}-1\right)}{2}K_{3}\left(\Sigma_{3}\cap\mathscr{A}_{3}^{\left(4\right)}\right)\right)\\ &\quad\times\prod_{p\mid c\left(\Sigma\right)}K_{p}\left(\Sigma_{p}\right)\left(1+b_{p}\right)\prod_{p\in\mathcal{P}\cup\left\{3\right\}}\left(1+{n_{p}}^{z}b_{p}\right).\end{aligned}$ (30)

### v2

@@ 1 Introduction
Our proof largely follows the strategy of McGown and Tucker [MT23]. The improved uniform estimates of [BTT24] allow us to obtain the improved error estimates in our main theorems.
@@
The organization of this paper is as follows: In Section 2, we summarize the necessary results on the density of cubic fields. In particular, we state the uniform error term estimates for counting cubic fields with prescribed local specifications, which are the results in [BTT24]. In Section 3, we provide the proofs of our main results, Theorems 1.3 and 1.4. We first recall Fröhlich’s formula, which shows that the genus number is determined by local properties. Then, using the discriminant decomposition we analyze the counting functions for cubic fields with fixed squarefree parts of the discriminant. By summing these contributions, we derive the asymptotic formulas for the distribution of genus numbers.
@@ 2 Preliminaries on Cubic Field Densities
In studying the distribution of cubic fields satisfying given local conditions, the following two theorems (Theorem 2.1 and Theorem 2.2) serve as our main tools. Theorems 2.1 and 2.2 are Theorems 1.3 and 1.4 in [BTT24], respectively.
Theorem 2.1 (Bhargava-Taniguchi-Thorne).
For an almost ordinary collection of local specifications $\Sigma=\left(\Sigma_{v}\right)_{v}$, let $N\left(X,\Sigma\right)$ denote the number of $F\in\Sigma$ satisfying $\left|\operatorname{Disc}\left(F\right)\right|\leq X$ Then, we have
$N\left(X,\Sigma\right)=C\left(\Sigma\right)\frac{1}{12\zeta\left(3\right)}X+K\left(\Sigma\right)\frac{4\zeta\left(1/3\right)}{5{\Gamma\left(2/3\right)}^{3}\zeta\left(5/3\right)}X^{5/6}+O_{\varepsilon}\left(E\left(\Sigma\right)X^{2/3+\varepsilon}\right)$ (7)
@@
where $v$ runs over all places of $\mathbb{Q}$. The local factors $C_{\infty}\left(\Sigma_{\infty}\right)$, $K_{\infty}\left(\Sigma_{\infty}\right)$, and $E_{\infty}\left(\Sigma_{\infty}\right)$ are defined by $C_{\infty}\left(\mathbb{R}^{3}\right)\coloneqq C^{+}$, $C_{\infty}\left(\mathbb{R}\times\mathbb{C}\right)\coloneqq C^{-}$, $K_{\infty}\left(\mathbb{R}^{3}\right)\coloneqq K^{+}$, $K_{\infty}\left(\mathbb{R}\times\mathbb{C}\right)\coloneqq K^{-}$, and $E_{\infty}\left(\Sigma_{\infty}\right)\coloneqq 1$. The local factors $C_{p}\left(\Sigma_{p}\right)$ and $K_{p}\left(\Sigma_{p}\right)$ for a prime $p$ are defined as in Theorem 1.3 of [BTT24].
@@
Theorem 2.2 (Bhargava-Taniguchi-Thorne).
For each collection of cubic local specifications $\Sigma=\left(\Sigma_{v}\right)_{v}$, let $E\left(X,\Sigma\right)$ denote the error term in estimating $N\left(X,\Sigma\right)$ in Theorem 2.1. Then for any $\varepsilon>0$, we have
$\sum_{r\leq R}\sum_{s}\left|E\left(X_{r,s},\Sigma^{r,s}\right)\right|\ll_{\varepsilon}X^{2/3+\varepsilon}R^{2/3}\prod_{p\mid U}E_{p}\left(\Sigma_{p}\right)$ (9)
@@ 3 Proofs of Main Results
The contribution of cyclic fields is negligible for the main results of this paper. First, regarding Theorem 1.3, the number of cyclic cubic fields with discriminant up to $X$ is known to be $O\left(X^{1/2}\right)$ (see [MT23, Section 2]). Second, regarding Theorem 1.4, we verify that the contribution of cyclic fields to the moments is negligible. Recall that for a squarefree positive integer $f$ coprime to $3$, the number of cyclic cubic fields $F$ satisfying $f_{F}=f$ is at most $2^{\omega\left(f\right)}$ (see [MT23, Section 4]), and their genus numbers satisfy $g_{F}\leq 3^{\omega\left(f\right)}$. Under the assumption $\operatorname{Re}\left(z\right)\leq G\left(X\right)$ with $G\left(X\right)=o\left(\log\log X\right)$, Théorème 11 of [Rob83] implies that for any $\varepsilon>0$, there exists a constant $C^{\prime}_{G,\varepsilon}$ depending on $G$ and $\varepsilon$ such that for all $X>C^{\prime}_{G,\varepsilon}$, we have
@@
Then we have $M\left(X,f\right)=N\left(X,\Sigma^{f}\right)$. For convenience, we define
@@
$\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E\left(X,f\right)\right|=\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E\left(X,\Sigma^{f}\right)\right|\ll_{\varepsilon}X^{2/3+\varepsilon}\prod_{p\mid 3c\left(\Sigma\right)t}E\left(\Sigma_{p}\right)\ll_{\varepsilon}{c\left(\Sigma\right)}^{2/3}X^{2/3+\varepsilon}.$ (21)
@@
Next, we consider $M^{\prime}\left(X,f\right)$. According to [TT14, Lemma 2.1], for a cubic field $F$,
[…]
where we note that the quotient $\operatorname{Disc}\left(F_{3}\right)/\operatorname{Disc}_{3}\left(F_{3}\right)$ is well-defined modulo $3$ because $u^{2}\equiv 1\pmod{3}$ for any $u\in\mathbb{Z}_{3}^{\times}$. Let $\mathscr{T}_{3}^{\left(1\right)}$ denote the subset of $\mathscr{A}_{3}$ consisting of cubic étale $\mathbb{Q}_{3}$-algebras $F_{3}$ such that $\operatorname{Disc}_{3}\left(F_{3}\right)=3^{4}$ and $\operatorname{Disc}\left(F_{3}\right)/\operatorname{Disc}_{3}\left(F_{3}\right)\equiv 1\pmod{3}$. According to [TT13, Section 6.2], $\mathscr{T}_{3}^{\left(1\right)}$ is the set of three totally ramified cubic étale $\mathbb{Q}_{3}$-algebras with the generating polynomials $x^{3}-3x^{2}+3u$ for $u\in\left\{1,4,7\right\}$.
@@
$\alpha^{\prime}\left(f\right)$ $\coloneqq\frac{\left(1+a_{3}\right)C_{\infty}\left(\Sigma_{\infty}\right)C_{3}\left(\Sigma_{3}\cap\mathscr{T}_{3}^{\left(1\right)}\right)}{12\zeta\left(2\right)}a_{tf}\prod_{p\mid c\left(\Sigma\right)}C_{p}\left(\Sigma_{p}\right)\left(1+a_{p}\right),$
$\beta^{\prime}\left(f\right)$ $\coloneqq\frac{4\left(1+b_{3}\right)\zeta\left(1/3\right)K_{\infty}\left(\Sigma_{\infty}\right)K_{3}\left(\Sigma_{3}\cap\mathscr{T}_{3}^{\left(1\right)}\right)}{5{\Gamma\left(2/3\right)}^{3}}b_{tf}\prod_{p}\left(1-\frac{p^{1/3}+1}{p\left(p+1\right)}\right)\prod_{p\mid c\left(\Sigma\right)}K_{p}\left(\Sigma_{p}\right)\left(1+b_{p}\right),$
@@
$\Sigma^{3,f}_{v}\coloneqq\begin{dcases}\Sigma_{v}\cap\mathscr{T}_{v}&\text{if }v\mid tf;\\ \Sigma_{v}\cap\mathscr{A}^{\prime}_{v}&\text{if }v\nmid 3tf;\\ \Sigma_{v}&\text{if }v=\infty;\\ \Sigma_{3}\cap\mathscr{T}_{3}^{\left(1\right)}&\text{if }v=3.\end{dcases}$ (23)
Then we have $M^{\prime}\left(X,f\right)=N\left(X,\Sigma^{3,f}\right)$. For all $v\neq 3$, we have $\Sigma^{3,f}_{v}=\Sigma^{f}_{v}$. Therefore, as in the previous proposition, $C\left(\Sigma^{3,f}\right)$, $K\left(\Sigma^{3,f}\right)$, and $E\left(\Sigma^{3,f}\right)$ are computed accordingly, and the proof of the first part is complete.
@@
$\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E^{\prime}\left(X,f\right)\right|=\sum_{f\in\operatorname{SF}\left(\Sigma\right)}\left|E\left(X,\Sigma^{3,f}\right)\right|\ll_{\varepsilon}c\left(\Sigma\right)^{2/3}X^{2/3+\varepsilon}.$ (24)
@@
$I_{3}\left(k,\Sigma\right)$ $=C_{3}\left(\Sigma_{3}\setminus\mathscr{T}_{3}^{\left(1\right)}\right)A_{k-l}\left(\Sigma\right)+C_{3}\left(\Sigma_{3}\cap\mathscr{T}_{3}^{\left(1\right)}\right)A_{k-l-1}\left(\Sigma\right),$
$J_{3}\left(k,\Sigma\right)$ $=K_{3}\left(\Sigma_{3}\setminus\mathscr{T}_{3}^{\left(1\right)}\right)B_{k-l}\left(\Sigma\right)+K_{3}\left(\Sigma_{3}\cap\mathscr{T}_{3}^{\left(1\right)}\right)B_{k-l-1}\left(\Sigma\right).$
@@
The last equality is justified because $M\left(X,f\right)=M^{\prime}\left(X,f\right)=0$ for $f\geq\sqrt{X}$. We introduce this truncation to bound the contribution of the weights $3^{\psi\left(f\right)z}$ in the tail of the series in the above derivation. Defining $\alpha^{\Sigma}\left(z\right)$ and $\beta^{\Sigma}\left(z\right)$ as
@@
$\alpha^{\Sigma}\left(z\right)$ $\coloneqq\begin{aligned} \frac{3^{lz}\cdot C_{\infty}\left(\Sigma_{\infty}\right)}{12\zeta\left(2\right)}a_{t}&\left(C_{3}\left(\Sigma_{3}\right)+\left(3^{z}-1\right)C_{3}\left(\Sigma_{3}\cap\mathscr{T}_{3}^{\left(1\right)}\right)\right)\\ &\quad\quad\quad\quad\quad\times\prod_{p\mid c\left(\Sigma\right)}C_{p}\left(\Sigma_{p}\right)\left(1+a_{p}\right)\prod_{p\in\mathcal{P}\cup\left\{3\right\}}\left(1+{n_{p}}^{z}a_{p}\right),\end{aligned}$ (28)
$\beta^{\Sigma}\left(z\right)$ $\coloneqq\begin{aligned} \frac{3^{lz}\cdot 4\zeta\left(1/3\right)K_{\infty}\left(\Sigma_{\infty}\right)}{5{\Gamma\left(2/3\right)}^{3}}b_{t}&\prod_{p}\left(1-\frac{p^{1/3}+1}{p\left(p+1\right)}\right)\left(K_{3}\left(\Sigma_{3}\right)+\left(3^{z}-1\right)K_{3}\left(\Sigma_{3}\cap\mathscr{T}_{3}^{\left(1\right)}\right)\right)\\ &\quad\times\prod_{p\mid c\left(\Sigma\right)}K_{p}\left(\Sigma_{p}\right)\left(1+b_{p}\right)\prod_{p\in\mathcal{P}\cup\left\{3\right\}}\left(1+{n_{p}}^{z}b_{p}\right).\end{aligned}$ (29)